Suppose in its 2017 annual report that McDonalds Corporation reports beginning total assets of $29.80 billion, ending total assets of $31.80 billion, net sales of $22.85 billion, and net income of $4.40 billion. (a) Compute McDonalds return on assets. (Round return on assets to 2 decimal places, e.g. 5.12%.) McDonalds return on assets (b) Compute McDonalds asset turnover. (Round asset turnover to 2 decimal places, e.g. 5.12.) McDonalds asset turnover times
Answer #1
a
Average total assets = (29.80+31.80)/2= $30.8 billion
Mcdonald's return on assets = Net income/Average total assets = 4.4/30.8= 14.29%
b
Mcdonald's asset turnover = Net sales/Average total assets = 22.85/30.8= 0.74 times